OpenStudy (anonymous):

Statistic, Definition of Independent 1) P(A|B) = P(A) 2) P(B|A) = P(B) 3) P(A Union B) = P(A)*P(B) explain why definition 3 is true for proving cases of independent

jimthompson5910 (jim_thompson5910):

are you asking if that is true or false?

jimthompson5910 (jim_thompson5910):

3) P(A Union B) = P(A)*P(B) is not true assuming A and B are independent

OpenStudy (anonymous):

hmmm, I got these definition off a textbook and is wondering why P(A)*P(B) resulting in proving the set function is independent. Further explanation please.

jimthompson5910 (jim_thompson5910):

The definition is really \[\Large P(A \cap B) = P(A)*P(B)\]

jimthompson5910 (jim_thompson5910):

Notice how I used an intersection symbol and NOT a union symbol

OpenStudy (anonymous):

Omg, you're right. What I meant is intersect!

jimthompson5910 (jim_thompson5910):

It turns out that \[\Large P(A \cap B) = P(A)*P(B|A)\] assuming A and B aren't independent. However, notice how that if A and B were independent, then P(B|A) = P(B) which leads us back to \[\Large P(A \cap B) = P(A)*P(B)\]

jimthompson5910 (jim_thompson5910):

you can do the same form of argument to go from \[\Large P(A \cap B) = P(A|B)*P(B)\] to \[\Large P(A \cap B) = P(A)*P(B)\] if A and B were independent